OpenStudy (anonymous):

Sarah’s dad tells her she can have the greater of the following two amounts for her allowance (in dollars): 2 times the difference of 5 and 1 5 times 1 minus 2 What is Sarah’s allowance, in dollars?

OpenStudy (mysesshou):

the difference of 5 and 1 means subtract 2 times that amount means multiply. so you get 2 * (5-1) = 2*4 = 8

OpenStudy (mysesshou):

as for the 5 times 1 minus 2, exactly how was that written?

OpenStudy (mysesshou):

times happens before minus, so then it would be (5*1) -2 = 4-2 = 2 maybe.

OpenStudy (chriss):

going by order of operations multiplication comes first so it would be 5(1)-2 so 5-2 or 3

OpenStudy (mysesshou):

* eh, yea, 5-2 = 3